E Co. purchased and placed in service the following assets in 2020:1. Computer (5-year property) purchased for $3,000,000 and placed in service on June 30, 2020;2. Tractor unit for use over the road (3-year property) purchased for $2,000,000 and placed in service on September 30, 2020;3. Equipment (7-year property) purchased for $5,000,000 and placed in service on October 1, 2020. All assets are depreciated under MACRS GDS (no SL election). E Co. is not eligible for179 expenses (phased out), and E Co. elected out of bonus depreciation. What is E Co.'s total depreciation deduction for 2020?
Answer #1

Computer :

Basis: $3,000,000
Date Placed In Service: July of 2020
Recovery Period: 5 Years
Depreciation Method: 200% Declining Balance Method
IRS Convention: Half-Year Convention

MACRS Depreciation: 3,000,000*20% = 600,000

Tractor:

Basis: $2,000,000
Date Placed In Service: October of 2020
Recovery Period: 3 Years
Depreciation Method: 200% Declining Balance Method
IRS Convention: Half-Year Convention

MACRS Depreciation: 2,000,000*33.33% = 666,667

Equipment:

Basis: $5,000,000
Date Placed In Service: November of 2020
Recovery Period: 7 Years
Depreciation Method: 200% Declining Balance Method
IRS Convention: Half-Year Convention

MACRS Depreciation: 5,000,000*14.29% = 714,286

Total depreciation to be recorded in 2020= $ 600,000 + $ 666,667 + $ 714,286

= $ 1,980,953